OpenStudy (anonymous):

Solve: 3x+(1-x)4= 2(4-x)

OpenStudy (anonymous):

3x+(1-x)4= 2(4-x) Are you sure that the (1-x)4 part is not (1-x)^4?

OpenStudy (anonymous):

yep i am sure

OpenStudy (anonymous):

3x+(1-x)4= 2(4-x) Can be rewritten as 3x + 4(1 - x) = 2(4 - x) This will come to 3x + 4 - 4x = 8 - 2x -> -x+4 = 8 - 2x -> -x = 4 - 2x -> x = 4

OpenStudy (anonymous):

thank you @Cosmichaotic

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

what/

OpenStudy (anonymous):

Nvm, that's right. Np.